OpenStudy (anonymous):

solve by factoring: 6-7u=3u² Please explain how you did this! Thanks!

OpenStudy (anonymous):

3u^2+7u-6=0

OpenStudy (anonymous):

take constant term and 1st term

OpenStudy (anonymous):

well I've already gotten that first part

OpenStudy (anonymous):

and find possible numbers that will go

OpenStudy (chaise):

Factor: 3u^2+7u-6=0 You need multiple of three and multiples of neg 6 which simultaneously multiply and add to give positive seven. (u+3)(3u-2)

OpenStudy (anonymous):

isn't -3u²+6-7u=0

OpenStudy (anonymous):

aorry thats wrong

OpenStudy (anonymous):

well I've already gotten that first part

OpenStudy (anonymous):

Which one mine? Yours?

OpenStudy (anonymous):

mine and yours this number cant be factored and urs the middle term is not right

OpenStudy (anonymous):

well I've already gotten that first part

OpenStudy (anonymous):

this isn't for pleasure, this a math problem for school! I don't really think they would give me something impossible to solve!

OpenStudy (anonymous):

nvm sorry chaise is right sorry im fasting